### 16 Dec 2023 v1

# Fractional Deep Reinforcement Learning for Age-Minimal Mobile Edge Computing

Abstract
Mobile edge computing (MEC) is a promising paradigm for real-time applications with intensive computational needs (e.g., autonomous driving), as it can reduce the processing delay. In this work, we focus on the timeliness of computational-intensive updates, measured by Age-of-Information (AoI), and study how to jointly optimize the task updating and offloading policies for AoI with fractional form. Specifically, we consider edge load dynamics and formulate a task scheduling problem to minimize the expected time-average AoI. The uncertain edge load dynamics, the nature of the fractional objective, and hybrid continuous-discrete action space (due to the joint optimization) make this problem challenging and existing approaches not directly applicable. To this end, we propose a fractional reinforcement learning (RL) framework and prove its convergence. We further design a model-free fractional deep RL (DRL) algorithm, where each device makes scheduling decisions with the hybrid action space without knowing the system dynamics and decisions of other devices. Experimental results show that our proposed algorithms reduce the average AoI by up to $57.6\%$ compared with several non-fractional benchmarks.
1 Introduction
Background and Motivations
The next-generation network demands mobile devices (e.g., smartphones and Internet-of-Things devices) to generate zillions of bytes of data and accomplish unprecedentedly computationally intensive tasks. Mobile devices, however, will be unable to timely process all their tasks locally due to their limited computational resources. To fulfill the low latency requirement, mobile edge computing (Mao et al. 2017) (MEC), also known as multi-access edge computing (Porambage et al. 2018), has become an emerging paradigm distributing computational tasks and services from the network core to the network edge. By enabling mobile devices to offload their computational tasks to nearby edge nodes, MEC can reduce the task processing delay.
On the other hand, the proliferation of real-time and computation-intensive applications (e.g., cyber-physical systems) has significantly boosted the demand for information freshness , e.g., (Yates et al. 2021; Kaul, Yates, and Gruteser 2012; Shisher and Sun 2022), in addition to low latency. For example, the real-time velocity and location knowledge of the surrounding vehicles is crucial in achieving safe and efficient autonomous driving. Another emerging example is metaverse applications, in which users anticipate real-time virtual reality services and real-time control over their avatars. In these applications, the experience of users depends on how fresh the received information is rather than how long it takes to receive that information. Such a requirement motivates a new network performance metric, namely Age of Information (AoI) (Yates et al. 2021; Kaul, Yates, and Gruteser 2012; Shisher and Sun 2022). It measures the time elapsed since the most up-to-date data (computational results) was received.
While the majority of existing studies on MEC were concerned about delay reduction (e.g., (Wang et al. 2021; Tang and Wong 2022)), most of real-time applications mentioned above concern about fresh status updates, while delay itself does not directly reflect timeliness. Here we highlight the huge difference between delay and AoI. Specifically, task delay takes into account only the duration between when the task is generated and when the task output has been received by the mobile device. Thus, under less frequent updates (i.e., when tasks are generated in a lower frequency), task delays are naturally smaller. This is because infrequent updates lead to empty queues and hence reduced queuing delays of the tasks. In contrast, AoI takes into account both the task delay and the freshness of the task output. Thus, to minimize the AoI with computational-intensive tasks, the update frequency needs to be neither too high nor too low in order to reduce the delay of each task while ensuring the freshness of the most up-to-date task output. [^1] Interesting readers may refer to (Talak and Modiano 2021)(Talak and Modiano 2021) for a more detailed analysis of tradeoffs between AoI and delay. More importantly, such a difference between delay and AoI leads to a counter-intuitive important phenomenon in designing age-minimal scheduling policy: upon the reception of each update, the mobile device may need to wait for a certain amount of time to generate the next new task (Sun et al. 2017).
Therefore, the age-minimal MEC systems necessitate meticulous design of a scheduling policy for each mobile device, which should encompass two fundamental decisions. The first decision is the updating decision, i.e., upon completion of a task, how long should a mobile device wait for generating the next one. The second is the task offloading decision, i.e., whether to offload the task or not? If yes, which edge node to choose? Although existing works on MEC have addressed the task offloading decision (e.g., (Tang and Wong 2022; Ma et al. 2022; Zhao et al. 2022; Zhu et al. 2022; He et al. 2022; Chen et al. 2022)) and some studies considered AoI (e.g., (Zhu et al. 2022; He et al. 2022; Chen et al. 2022)), they did not consider designing the task updating policy to improve data timeliness.
In this paper, we aim to answer the following question:
Key Question.
How should mobile devices optimize their updating and offloading policies of hybrid action space in dynamic MEC systems in order to minimize their fractional objectives of AoI?
Solution Approach and Contributions
In this work, we take into account system dynamics in MEC systems and aim at designing distributed AoI-minimal DRL algorithms to jointly optimize task updating and offloading. We first propose a novel fractional RL framework, incorporating reinforcement learning techniques and Dinkelbach’s approach (for fractional programming) in (Dinkelbach 1967). We further propose a fractional Q-Learning algorithm and analyze its convergence. To address the hybrid action space, we further design a fractional DRL-based algorithm. Our main contributions are summarized as follows:
  • Joint Task Updating and Offloading Problem: We formulate the joint task updating and offloading problem that takes into account unknown system dynamics. To the best of our knowledge, this is the first work designing the joint updating and offloading policy for age-minimal MEC.
  • Fractional RL Framework: To overcome fractional objective of the average AoI, we propose a novel fractional RL framework. We further propose a fractional Q-Learning algorithm. We design a stopping condition, leading to a provable linear convergence rate without the need of increasing inner-loop steps.
  • Fractional DRL Algorithm: We overcome unknown dynamics and hybrid action space of offloading and updating decisions and propose a fractional DRL-based distributed scheduling algorithm for age-minimal MEC, which extends the dueling double deep Q-network (D3QN) and deep deterministic policy gradient (DDPG) techniques into our proposed fractional RL framework.
  • Performance Evaluation: Our algorithm significantly outperforms the benchmarks that neglect the fractional nature with an average AoI reduction by up to $57.6\%$. In addition, the joint optimization of offloading and updating can further reduce the AoI by up to $31.3\%$.
2 Literature Review
Mobile Edge Computing: Existing excellent works have conducted various research questions in MEC, including resource allocation (e.g., (Wang et al. 2022b)), service placement (e.g., (Taka, He, and Oki 2022)), and proactive caching (e.g., (Liu et al. 2022a)). Task offloading (Wang, Ye, and Lui 2022; Ma et al. 2022; Chen and Xie 2022), as another main research question in MEC, attracting considerable attention. To address the unknown system dynamics and reduce task delay, many existing works have proposed DRL-based approaches to optimize the task offloading in a centralized manner (e.g., (Huang, Bi, and Zhang 2020; Tuli et al. 2022)). As in our work, some existing works have proposed distributed DRL-based algorithms (e.g., (Tang and Wong 2022; Liu et al. 2022b; Zhao et al. 2022)) which do not require the global information. Despite the success of these works in reducing the task delay, these approaches are NOT easily applicable to age-minimal MEC due to the aforementioned challenges of fractional objective and hybrid action space.
Age of Information: Kaul et al. first introduced AoI in (Kaul, Yates, and Gruteser 2012). Assuming complete and known statistical information, the majority of this line of work mainly focused on the optimization and analysis of AoI in queueing systems and wireless networks (see (Zou, Ozel, and Subramaniam 2021; Chiariotti et al. 2021; Kuang et al. 2020), and a survey in (Yates et al. 2021)). Zou et al. in (Zou, Ozel, and Subramaniam 2021), Chiariotti et al. in (Chiariotti et al. 2021), and Kuang et al. in (Kuang et al. 2020). The above studies analyzed simple single-device-single-server models and hence did not consider offloading.
A few studies investigated DRL algorithm design to minimize AoI in various application scenarios, including wireless networks (e.g., (Ceran, Gündüz, and György 2021)), Internet-of-Things (e.g., (Akbari et al. 2021; Wang et al. 2022a)), vehicular networks (e.g., (Chen et al. 2020)), and UAV-aided networks (e.g., (Hu et al. 2020; Wu et al. 2021)). This line of work mainly focused on optimal resource allocation and trajectory design. Existing works considered AoI as the performance metric for task offloading in MEC and proposed DRL-based approaches to address the AoI minimization problem. Chen et al. in (Chen et al. 2022) considered AoI to capture the freshness of computation outcomes and proposed a multi-agent DRL algorithm. However, these works focused on designing task offloading policy but did not optimize updating policy. Most importantly, all aforementioned approaches did not account for fractional RL and hence cannot directly tackle our considered problem.
RL with Fractional Objectives: Research on RL with fractional objectives is currently limited. Ren et. al. introduced fractional MDP (Ren and Krogh 2005). Reference (Tanaka 2017) further studied partially observed MDPs with fractional rewards. However, RL methods were not considered in these studies. Suttle et al. (Suttle et al. 2021) proposed a two-timescale RL algorithm for the fractional cost, but it requires additional fixed reference states in the Q-learning update process to approximate the outer loop update and leave finite-time convergence analysis unsettled.
3 System Model
Consider $M$ mobile devices and $N$ edge nodes, which are in set =$\mathcal{M}=\{1,2,\ldots,M\}$ and set $\mathcal{N}=\{1,2,\ldots N\}$, respectively. We consider an infinite-horizon continuous-time system model illustrated in Fig. 1.
Figure 1: An illustration of an MEC system with a mobile device $m\in\mathcal{M}$ and an edge node $n\in\mathcal{N}$ wherethe tasks offloaded by different mobile devices are represented using different colors.
Device Model
Each mobile device $m\in\mathcal{M}$ has a task generator, a scheduler, and a monitor. The task generator generates new tasks for processing while the scheduler determines where to process the tasks and the task output is sent to the monitor. We refer to a task output received by the monitor as an update.
Task Generator: We consider a generate-at-will model, as in (Sun et al. 2017), i.e., each task generator can decide when to generate the next new task. At the time when task $k-1$ of mobile device $m\in\mathcal{M}$ has been completed (denoted by time $t^{\prime}_{m,k-1}$), the task generator observes the task delay $Y_{m,k-1}$ and makes a decision on $Z_{m,k}$, i.e., the waiting time for generating the next task $k$. Let $\bm{s}^{\textsc{u}}_{m}(k)$ and $\bm{a}^{\textsc{u}}_{m}(k)$ denote the state and action with task $k$ of mobile device $m$ respectively: $\bm{s}^{\textsc{u}}_{m}(k)=Y_{m,k-1},~{}\bm{a}^{\textsc{u}}_{m}(k)=Z_{m,k},~{}k\in\mathcal{K},m\in\mathcal{M}$.
Let $\mathcal{S}^{\textsc{u}}=(0,\bar{Y}]$ and $\mathcal{A}^{\textsc{u}}\in[0,\bar{Z}]$ denote the state and action space, respectively. Let ${\pi}^{\textsc{u}}_{m}:\mathcal{S}^{\textsc{u}}\rightarrow\mathcal{A}^{\textsc{u}}$ denote the updating policy of mobile device $m\in\mathcal{M}$ that maps from $\mathcal{S}^{\textsc{u}}$ to $\mathcal{A}^{\textsc{u}}$. Specifically, let $t_{m,k}$ denote the time stamp when the task generator of mobile device $m$ generates task $k$, after which the task is sent to the scheduler. The transmission time is considered negligible and $t_{m,k+1}=t^{\prime}_{m,k}+Z_{m,k+1}$. From (Sun et al. 2017), the optimal waiting strategy may outperform the zero-wait policy, i.e., $Z_{m,k}$ may not necessarily be zero and requires proper optimization.
Scheduler: At the time when task $k$ of mobile device $m\in\mathcal{M}$ is generated (i.e., time $t_{m,k}$), the task scheduler observes the queue lengths of edge nodes and makes the offloading decision denoted by $x_{m,k}\in\{0\}\cup\mathcal{N}$. Let $\bm{s}^{\textsc{o}}_{m}(k)$ denote the state vector associated with task $k$ of mobile device $m$: [^2] Under the system model in Section 3, observing the queue lengths of edge nodes is sufficient for mobile devices to learn their offloading policies. Under a more complicated system (e.g., with device mobility), mobile devices may need to observe additional state information, while our proposed fractional RL framework and DRL algorithm are still applicable with the extended state vector. $\bm{s}^{\textsc{o}}_{m}(k)=\bm{q}(t_{m,k}),~{}k\in\mathcal{K},m\in\mathcal{M}$, where $\bm{q}(t_{m,k})=(q_{n}(t_{m,k}),n\in\mathcal{N})$ corresponds to the queue lengths of all edge nodes. We assume that edge nodes send their queue length information upon the requests of mobile devices. Since a generator generates a new task only after the previous task has been processed, the queue length is less than or equal to $M$. Thus, it can be encoded in $O(\log_{2}M)$ bits, which incurs only small signaling overheads. Let $\mathcal{S}^{\textsc{o}}=\mathcal{M}^{1\times N}$ denote the state space. Let $\bm{a}^{\textsc{o}}_{m}(k)$ denote the action associated with task $k$ of mobile device $m$. Thus, $\bm{a}^{\textsc{o}}_{m}(k)=x_{m,k},~{}~{}k\in\mathcal{K},m\in\mathcal{M}$. Let $\mathcal{A}^{\textsc{o}}\in\{0\}\cup\mathcal{N}$ denote the offloading action space. Let ${\pi}^{\textsc{o}}_{m}$ denote the task offloading policy of mobile device $m\in\mathcal{M}$.
If mobile device $m$ processes task $k$ locally, then let $\tau^{\rm{local}}_{m,k}$ (in seconds) denote the service time of mobile device $m\in\mathcal{M}$ for processing task $k$. The value of $\tau^{\rm{local}}_{m,k}$ depends on the size of task $k$ and the real-time processing capacity of mobile device $m$ (e.g., whether the device is busy in processing tasks of other applications), which are unknown a priori. If mobile device $m$ offloads task $k$ to edge node $n\in\mathcal{N}$, then let $\tau^{\rm{tran}}_{n,m,k}$ (in seconds) denote the service time of mobile device $m\in\mathcal{M}$ for sending task $k$ to edge node $n$. The value of $\tau^{\rm{tran}}_{n,m,k}$ depends on the time-varying wireless channels and is unknown a priori. We assume that $\tau^{\rm{local}}_{m,k}$ and $\tau^{\rm{tran}}_{n,m,k}$ are random variables that follow exponential distribution (Tang et al. 2021; Liu et al. 2022b; Zhu and Gong 2022), respectively, which are independent of edge loads.
Edge Node Model
Upon receiving a task offloaded by a mobile device, edge node $n\in\mathcal{N}$ enqueues the task for processing. The queue may store the tasks offloaded by multiple mobile devices, as shown in Fig. 1. Suppose the queue operates in a first-in-first-out (FIFO) manner (Yates et al. 2021).
Let $w^{\rm{edge}}_{n,m,k}$ (in seconds) denote the time duration that task $k$ of mobile device $m\in\mathcal{M}$ waits at the queue of edge node $n$. Let $\tau^{\rm{edge}}_{n,m,k}$ (in seconds) denote the service time of edge node $n$ for processing task $k$ of mobile device $m$. The value of $\tau^{\rm{edge}}_{n,m,k}$ depends on the size of task $k$ and may be unknown a priori. We assume that $\tau^{\rm{edge}}_{n,m,k}$ is a random variable following exponential distribution as well. In addition, the value of $w^{\rm{edge}}_{n,m,k}$ depends on the processing time of the tasks placed in the queue (of edge node $n$) ahead of the task $k$ of device $m$, where those tasks are possibly offloaded by mobile devices other than device $m$. Thus, since mobile device $m$ does not know the offloading behaviors of other mobile devices a priori, it does not know the value of $w^{\rm{edge}}_{n,m,k}$ beforehand.
Age of Information
The age of information (AoI) for mobile device $m$ at time stamp $t$ (Yates et al. 2021) is given by
$\Delta_{m}(t)=t-U_{m}(t),~{}~{}\forall m\in\mathcal{M},t\geq 0,$ (1)
where $U_{m}(t)\triangleq\max_{k}[t_{m,k}|t^{\prime}_{m,k}\leq t]$ stands for the time stamp of the most recently completed task.
We use $Y_{m,k}\triangleq t_{m,k}^{\prime}-t_{m,k}$ to denote the delay of task $k$, i.e., the time it takes to complete task $k$. Thus,
$Y_{m,k}=\begin{cases}\tau^{\rm{local}}_{m,k},&x_{m,k}=0,\\ \tau^{\rm{tran}}_{n,m,k}+w^{\rm{edge}}_{n,m,k}+\tau^{\rm{tran}}_{n,m,k},&x_{m,k}=n\in\mathcal{N}.\end{cases}$ (2)
We consider a drop time $\bar{Y}$ (in seconds). That is, we assume that if a task has not been completely processed within $\bar{Y}$ seconds, the task will be dropped (Tang and Wong 2022; Li, Zhou, and Chen 2020). Meantime, the AoI keeps increasing until the next task is completed.
To capture the overall performance of mobile device $m$, we define the trapezoid area associated with time interval $[t_{m,k},t_{m,k+1})$ (Yates et al. 2021):
$A(Y_{m,k},Z_{m,k+1},Y_{m,k+1})\triangleq~{}$ $\frac{1}{2}(Y_{m,k}+Z_{m,k+1}+Y_{m,k+1})^{2}$
$-\frac{1}{2}Y_{m,k+1}^{2}.$ (3)
Based on (3), we can characterize the objective of mobile device $m$, i.e., to minimize the time-average AoI of each device $m\in\mathcal{M}$: (Yates et al. 2021)
${\Delta}^{(ave)}_{m}\triangleq$ $\liminf_{M\rightarrow\infty}\frac{\sum_{m=1}^{M}A(Y_{m,k},Z_{m,k+1},Y_{m,k+1})}{\sum_{m=1}^{M}(Y_{m,k}+Z_{m,k+1})}.$ (4)
Problem Formulation
Let $\bm{\pi}_{m}=({\pi}^{\textsc{u}}_{m},{\pi}^{\textsc{o}}_{m})$ denote the policy of mobile device $m\in\mathcal{M}$. This is a stationary policy that contains the mapping from $\mathcal{S}^{\textsc{u}}\times\mathcal{S}^{\textsc{o}}$ to $\mathcal{A}^{\textsc{u}}\times\mathcal{A}^{\textsc{o}}$. Given a stationary policy $\bm{\pi}_{m}$, the expected time-average AoI of mobile device $m\in\mathcal{M}$ is
$\mathbb{E}[\Delta_{m}^{(ave)}|\bm{\pi}_{m}]\triangleq\frac{\mathbb{E}[A(Y_{m,k},Z_{m,k+1},Y_{m,k+1})|\bm{\pi}_{m}]}{\mathbb{E}[Y_{m,k}+Z_{m,k+1}|\bm{\pi}_{m}]}.$ (5)
We take the expectation $\mathbb{E}[\cdot]$ over policy $\bm{\pi}_{m}$ and the time-varying system parameters, e.g., the time-varying processing duration as well as the edge load dynamics.
We aim at the optimal policy $\bm{\pi}_{m}^{*}$ for each mobile device $m\in\mathcal{M}$ to minimize its expected time-average AoI, i.e., [^3] If the AoI minimization problem is modeled as a game or multi-agent RL-based approaches (e.g., (Chen et al. 2022)), the policies of mobile devices will converge to Nash equilibrium rather tan reduce the AoI of each device, and they need additional signaling between mobile devices for state information sharing. Thus, in this work, to minimize time-average AoI, we model the policy design problem of each mobile device as an optimization problem (rather than game). The simulation results in Section 6 show that with our proposed approach, the scheduling policies of mobile devices can gradually converge.
$\bm{\pi}_{m}^{*}=\arg\mathop{\text{minimize}}_{\bm{\pi}_{m}}\quad$ $\mathbb{E}\left[\Delta_{m}^{(ave)}|\bm{\pi}_{m}\right].$ (6)
The fractional objective in (5) introduces a major challenge in designing the optimal policy, which is significantly different from conventional RL and DRL algorithms. Specifically, the difficulty of directly expressing the immediate reward (cost) of each action for the fractional RL problem. Specifically, it seems to be straightforward to define the reward (or cost) function as either the instant AoI (i.e., $\Delta_{m}(t_{m,k}^{\prime})$) (Chen et al. 2022; He et al. 2022) or the average AoI during certain time interval (e.g., $[t_{m,k},t_{m,k+1})$). However, consider the time-average over infinite time horizon, neither minimizing $\mathbb{E}[\Delta_{m}(t_{m,k}^{\prime})|\bm{\pi}_{m}]$ nor $\mathbb{E}[A(Y_{m,k},Z_{m,k+1},Y_{m,k+1})/(Y_{m,k}+Z_{m,k+1})|\bm{\pi}_{m}]$ is equivalent to minimizing (5).
4 Fractional RL Framework
In this section, we propose a fractional RL framework for solving Problem (6).We first present a two-step reformulation of Problem (6). We then introduce the fractional RL framework, under which we present a fractional Q-Learning algorithm with provable convergence guarantees.
Dinkelbach’s Reformulation
With the proposed Problem 6 we consider the Dinkelbach’s reformulation and a discounted reformulation in the following. we define a reformulated AoI in an average-cost fashion:
$\mathbb{E}[\Delta_{m}^{\prime}|\bm{\pi}_{m},\gamma]\!\triangleq$ $\!\!\lim_{K\rightarrow\infty}\!\frac{1}{K}\!\sum_{k=1}^{K}\!\left\{\mathbb{E}[A(Y_{m,k},Z_{m,k+1},Y_{m,k+1})|\bm{\pi}_{m}]\right.$
$\left.-\gamma\mathbb{E}[Y_{m,k}+Z_{m,k+1}|\bm{\pi}_{m}]\right\}.$ (7)
Let $\gamma^{*}$ be the optimal value of Problem (6). Leveraging Dinkelbach’s method (Dinkelbach 1967), we have the following reformulated problem:
Lemma 1 ((Dinkelbach 1967)).
Problem (6) is equivalent to the following reformulated problem:
$\bm{\pi}_{m}^{*}=\arg\mathop{\rm{minimize}}_{\bm{\pi}_{m}}~{}\mathbb{E}[\Delta_{m}^{\prime}|\bm{\pi}_{m},\gamma^{*}],\quad\forall m\in\mathcal{M},$ (8)
where $\bm{\pi}_{m}^{*}$ is the optimal solution to Problem (6).
Since $\mathbb{E}[\Delta_{m}^{\prime}|\bm{\pi}_{m},\gamma^{*}]\geq 0$ for any $\pi$ and $\mathbb{E}[\Delta_{m}^{\prime}|\bm{\pi}_{m}^{*},\gamma^{*}]=0$, $\bm{\pi}_{m}^{*}$ is also optimal to the Dinkelbach reformulation. This implies the reformulation equivalence is also established for our stationary policy space.
Discounted Reformulation
Following Dinkelbach’s reformulation, we reformulate the problem in (8) one step further by considering a discounted objective. Let $\delta\in(0,1]$ be the discount factor, capturing how the objective is discounted in the future. We define
$\mathbb{E}[\Delta_{m}^{\delta}|\bm{\pi}_{m},\gamma]\triangleq$ $\sum_{k=1}^{\infty}\delta^{k}\left\{\mathbb{E}[A(Y_{m,k},Z_{m,k+1},Y_{m,k+1})|\bm{\pi}_{m}]\right.$
$\left.-\gamma\mathbb{E}[Y_{m,k}+Z_{m,k+1}|\bm{\pi}_{m}]\right\},\forall\gamma\geq 0.$ (9)
From (Puterman 2014), we can establish the asymptotic equivalence between the average formulation and the discounted formulation:
Lemma 2 (Asymptotic Equivalence (Puterman 2014)).
Given the optimal quotient value $\gamma^{*}$, Problems (6) and (8) are asymptotically equivalent to reformulation as follows:
$\bm{\pi}_{m}^{*}=\arg\mathop{\rm{minimize}}_{\bm{\pi}_{m}}~{}\lim_{\delta\rightarrow 1}\mathbb{E}[\Delta_{m}^{\delta}|\bm{\pi}_{m},\gamma^{*}],$ (10)
for all $m\in\mathcal{M}$, where $\bm{\pi}_{m}^{*}$ is the optimal solution to (8).
Therefore, the discounted reformulation in (10) serves as a good approximation of (7) when $\delta$ approaches $1$. The main purpose of introducing the discounted reformulation in (10) is that the majority of existing studies on RL (and DRL) algorithm design have focused on the discounted formulation (e.g., (Tang and Wong 2022; Zhao et al. 2022; Nguyen et al. 2021; Chen et al. 2022)). Hence, such an approximation provides us with a convention of designing new DRL algorithms for fractional MDP problems based on existing well-established DRL algorithms. We will therefore stick to the discounted reformulation for the rest of this paper.
Fractional MDP
We study the following general fractional MDP framework and drop index $m$ for the rest of this section.
Definition 1 (Fractional MDP).
A fractional MDP is defined as $(\mathcal{S},\mathcal{A},P,c_{N},c_{D},\delta)$, where $\mathcal{S}$ and $\mathcal{A}$ are the finite sets of states and actions, respectively; $P$ is the transition distribution; $c_{N}$ and $c_{D}$ are the cost functions, [^4] Since we aim at minimizing the time-average AoI, we consider minimizing long-term expected cost in this work. The proposed approach can be easily extended to the fraction MDP that maximizes long-term expected reward. and $\delta$ is a discount factor. We use $\mathcal{Z}$ to denote the joint state-action space, i.e., $\mathcal{Z}\triangleq\mathcal{S}\times\mathcal{A}$.
From Definition 1 and Lemmas 1 and 2, we have that Problem (6) has the equivalent Dinkelbach’s reformulation:
${\bm{\pi}}^{*}=\arg\mathop{\text{minimize}}_{{\bm{\pi}}}\lim_{K\rightarrow\infty}\mathbb{E}\left[\left.\sum_{k=1}^{K}\delta^{k}(c_{N}-\gamma^{*}c_{D})\right|{\bm{\pi}}\right],$ (11)
where we can see from Lemmas 1 and 2 that $\gamma^{*}$ satisfies [^5] Note that $\gamma^{*}$ always exists due to the minimization nature.
$\gamma^{*}=\mathop{\text{minimize}}_{{\bm{\pi}}}\lim_{K\rightarrow\infty}\frac{\left.\mathbb{E}\left[\sum_{k=0}^{K}\delta^{k}c_{N}\right|\bm{\pi}\right]}{\mathbb{E}\left[\left.\sum_{k=0}^{K}\delta^{k}c_{D}\right|\bm{\pi}\right]},$ (12)
Note that Problem (11) is a classical MDP problem, including an immediate cost, given by $c_{N}(\bm{s},\bm{a})-\gamma^{*}c_{D}(\bm{s},\bm{a})$. Thus, we can then apply a traditional RL algorithm to solve such a reformulated problem, such as Q-Learning or its variants (e.g., SQL in (Ghavamzadeh et al. 2011)).
However, the optimal quotient coefficient $\gamma^{*}$ and the transition distribution $P$ are unknown a priori. Therefore, one needs to design an algorithm that combines both fractional programming and RL algorithms to solve Problem (11) for a given $\gamma$ and seek the value of $\gamma^{*}$. To achieve this, we start by introducing the following definitions: Given a quotient coefficient $\gamma$, the optimal Q-function is
$Q^{*}_{\gamma}(\bm{s},\bm{a})\triangleq\min_{\bm{\pi}}Q^{\bm{\pi}}_{\gamma}(\bm{s},\bm{a}),~{}\forall(\bm{s},\bm{a})\in\mathcal{Z},$ (13)
where $Q^{\bm{\pi}}_{\gamma}(\bm{s},\bm{a})$ is the action-state function that satisfies the following Bellman’s equation: for all $(\bm{s},\bm{a})\in\mathcal{Z}$,
$Q^{\bm{\pi}}_{\gamma}(\bm{s},\bm{a})\triangleq c_{N}(\bm{s},\bm{a})-\gamma c_{D}(\bm{s},\bm{a})+\delta\mathbb{E}[Q^{\bm{\pi}}_{\gamma}(\bm{s},\bm{a})|\bm{\pi}].$ (14)
In addition, we can further decompose the optimal Q-function in (13) into the following two parts: $Q^{*}_{\gamma}(\bm{s},\bm{a})=N_{\gamma}(\bm{s},\bm{a})-\gamma D_{\gamma}(\bm{s},\bm{a})$ and, for all $(\bm{s},\bm{a})\in\mathcal{Z}$,
$N_{\gamma}(\bm{s},\bm{a})$ $=c_{N}(\bm{s},\bm{a})+\delta\mathbb{E}[N_{\gamma}(\bm{s},\bm{a})|\bm{\pi}^{*}],$ (15)
$D_{\gamma}(\bm{s},\bm{a})$ $=c_{D}(\bm{s},\bm{a})+\delta\mathbb{E}[D_{\gamma}(\bm{s},\bm{a})|\bm{\pi}^{*}].$ (16)
Algorithm 1 Fractional Q-Learning (FQL)
1:  for $k=1,2,\ldots,K$ do
2:     Initialize $\bm{s}_{m}(1)$
3:     for time slot $t\in\mathcal{T}$ do
4:        Observe the next state $\bm{s}_{m}(t+1)$
5:        Observe a set of costs $\{c_{m}(t^{\prime}),~{}t^{\prime}\in\widetilde{\mathcal{T}}_{m,t}\}$
6:        for each task $k_{m}(t^{\prime})$ with $t^{\prime}\in\widetilde{\mathcal{T}}_{m,t}$ do
7:           Send $(\bm{s}_{m}(t^{\prime}),\bm{a}_{m}(t^{\prime}),c_{m}(t^{\prime}),\bm{s}_{m}(t^{\prime}+1))$ to $n_{m}$
8:        end for
9:     end for
10:     $\gamma^{(k+1)}={N_{\gamma^{k}}(s,a_{k})}/{D_{\gamma^{k}}(s,a_{k})},$ where $a_{k}=\arg\max_{a}Q_{\gamma}^{T}(s,a)$.
11:  end for
Fractional Q-Learning Algorithm
In this subsection, we present a Fractional Q-Learning (FQL) algorithm in Algorithm 1, consisting of an inner loop with $E$ episodes and an outer loop. The key idea is to approximate the Q-function $Q_{\gamma_{i}}^{*}$ by $Q_{i}$ and then iterate $\{\gamma_{i}\}$.
One of the key innovations in Algorithm 1 is the design of the stopping condition, leading to the shrinking values of the uniform approximation errors of $Q_{i}$. This facilitates us to adapt the convergence proof in (Dinkelbach 1967) to our setting and prove the linear convergence rate of $\{\gamma_{i}\}$ without increasing the inner-loop time complexity.
We describe the details of the inner loop and the outer loop procedures of Algorithm 1 in the following:
  • Inner loop: For each episode $i$, given a quotient coefficient $\gamma_{i}$, we perform an (arbitrary) Q-Learning algorithm (as the Speedy Q-Learning in (Ghavamzadeh et al. 2011)) to approximate function $Q_{\gamma_{i}}^{*}(\bm{s},\bm{a})$ by $Q_{i}(\bm{s},\bm{a})$. Let $\bm{s}_{0}$ denote the initial state of any arbitrary episode, and $\bm{a}_{i}\triangleq\arg\max_{\bm{a}}Q_{i}(\bm{s}_{0},\bm{a})$ for all $i\in[E]\triangleq\{1,...,E\}$. We consider a stopping condition
    $\epsilon_{i}<-\alpha Q_{i}(\bm{s}_{0},\bm{a}_{i}),~{}\forall i\in[E],$ (17)
[…]
    $Q_{i}(\bm{s},\bm{a})=N_{i}(\bm{s},\bm{a})-\gamma_{i}D_{i}(\bm{s},\bm{a}).$ (18)
  • the outer loop to update the quotient coefficient:
    $\gamma_{i+1}=\frac{N_{i}(\bm{s}_{0},\bm{a}_{i})}{D_{i}(\bm{s}_{0},\bm{a}_{i})},\quad\forall i\in[E],$ (19)
    which will be shown to converge to the optimal value $\gamma^{*}$.
Convergence Analysis
We are ready to present the key convergence results of our proposed FQL algorithm (Algorithm 1) as follows.
Convergence of the outer loop
We start with analyzing the convergence of the outer loop:
Theorem 1 (Linear Convergence of Fractional Q-Learning).
If we select $\{T_{i}\}$ such that the uniform approximation error $\lVert Q^{*}_{\gamma_{i}}-Q_{i}\rVert\leq\epsilon_{i}$ holds with $\epsilon_{i}<-\alpha Q_{i}(\bm{s}_{0},\bm{a}_{i})$ for some $\alpha\in(0,1)$ and for all $i\in[E]$, then the sequence $\{\gamma_{i}\}$ generated by Algorithm 1 satisfies
$\frac{\gamma_{i+1}-\gamma^{*}}{\gamma_{i}-\gamma^{*}}\in(0,1),~{}\forall i\in[E]\text{ and }\lim_{i\rightarrow\infty}\frac{\gamma_{i+1}-\gamma^{*}}{\gamma_{i}-\gamma^{*}}=\alpha.$ (20)
That is, $\{\gamma_{i}\}$ converges to $\gamma^{*}$ linearly.
While the convergence proof in (Dinkelbach 1967) requires to obtain the exact solution in each episode, Theorem 1 generalizes this result to the case where we only obtain an approximated (inexact) solution in each episode. In addition to the proof techniques in (Dinkelbach 1967) and (Ghavamzadeh et al. 2011), our proof techniques include induction and exploiting the convexity of $Q_{i}(\bm{s},\bm{a})$. We present a proof sketch of Theorem 1 in Appendix A.
The significance of Theorem 1 is two-fold. First, Theorem 1 shows that Algorithm 1 achieves a linear convergence rate, even though it only attains an approximation of $Q_{\gamma}^{*}(\bm{s},\bm{a})$ Second, (17) is a well-behaved stopping condition.
Time Complexity of the inner loop
Although as $\{Q_{i}(\bm{s}_{0},\bm{a}_{i})\}$ is convergent to 00 and hence $\epsilon_{i}<-\alpha Q_{i}$ is getting more restrictive as $i$ increases, the steps needed $T_{i}$ in Algorithm 1 keep to be finite without increasing over episode $i$ based on (Ghavamzadeh et al. 2011). See Appendix B in detail.
Figure 2: Illustration on the proposed fractional DRL framework.
5 Fractional DRL Algorithm
In this section, we present a fractional DRL-based algorithm to approximate the Q-function in FQL algorithm and solve Problem (6) with DDPG (for continuous action space) (Lillicrap et al. 2015) and D3QN (for discrete action space) (Mnih et al. 2015) techniques [^6] Twin Delayed DDPG (TD3) (Fujimoto, van Hoof, and Meger 2018) and Proximal policy optimization (PPO) (Schulman et al. 2017) are also applicable to continuous action. Soft Actor-Critic for Discrete Action (SAC-discrete) (Christodoulou 2019) also works for discrete action space. We empirically show via extensive experiments that in our algorithm outperforms the other aforementioned techniques. Note that our proposed algorithm can be extended by incorporating other techniques. for the task updating and offloading processes in the decentralized manner, which is illustrated in Fig. 2. Appendix C shows detailed settings of networks with the reference of settings in (Tang and Wong 2022, Table I). Moreover, we design a cost function based on our fractional RL framework to ensure the convergence.
Cost Module
As in the proposed fractional RL framework, we consider a set of episodes $i\in[E]$ and introduce a quotient coefficient $\gamma_{i}$ for episode $i$. Consider mobile device $m$. Let $\bm{a}^{\textsc{h}}_{m,i}(k)\triangleq\{(\bm{a}^{\textsc{u}}_{m,i}(l),\bm{a}^{\textsc{o}}_{m,i}(l))\}_{l=1}^{k}$ denote the set of updating and offloading actions of mobile device $m\in\mathcal{M}$ made until task $k$ in episode $i$, where “H” refers to “history”. Recall that $\bm{s}_{0}$ is the initial state of any arbitrary episode. We define $N_{i}(\bm{s}_{0},\bm{a}^{\textsc{h}}_{m,i})$ and $D_{i}(\bm{s}_{0},\bm{a}^{\textsc{h}}_{m,i})$ as follows:
$N_{i}(\bm{s}_{0},\bm{a}^{\textsc{h}}_{m,i}(k))$ $=\sum_{l=1}^{k}A(Y_{m,l}^{i},Z_{m,l+1}^{i},Y_{m,l+1}^{i}),$ (21)
$D_{i}(\bm{s}_{0},\bm{a}^{\textsc{h}}_{m,i}(k))$ $=\sum_{l=1}^{k}({Y_{m,l}^{i}}+Z_{m,l+1}^{i}),$ (22)
where $Y_{m,l}^{i}$ and $Z_{m,l+1}^{i}$ are the delay of task $l$ and the time that the task generator waits for generating the next task $l+1$, respectively, for mobile device $m$. Note that $Y_{m,l}^{i}$ is a function of $\bm{a}^{\textsc{h}}_{m,i}(l)$, and $Z_{m,l+1}^{i}$ is a function of $\bm{a}^{\textsc{h}}_{m,i}(l)$ and $\bm{a}^{\textsc{u}}_{m,i}(l+1)$. The cost module keeps track of $\bm{a}^{\textsc{h}}_{m,i}(k)$ or equivalently $Y^{i}_{m,l}$ and $Z^{i}_{m,l}$ for all $l=1,...,k$ across the training process.
In step $k$ of episode $i$, [^7] Here, step $k$ refers to the updating and offloading processes as well as the training process associated with task $k$. a cost is determined and sent to the DDPG and D3QN modules. This process corresponds to the inner loop of the proposed fractional RL framework and is defined based on (21): for all $i\in[E],m\in\mathcal{M},k\in\mathcal{K},$
$c_{m,i}(k)\!=\!A(Y_{m,k}^{i},Z_{m,k+1}^{i},$ $Y_{m,k+1}^{i})\!-\!\gamma_{m,i}\!\cdot\!(Y_{m,k}^{i}\!+\!Z_{m,k+1}^{i}),$ (23)
where $A(Y_{m,k}^{i},Z_{m,k+1}^{i},Y_{m,k+1}^{i})$ stands for the area of a trapezoid in (3). The cost in (23) corresponds to an (immediate) cost function as in the fractional MDP problem in (11).
Finally, at the end of each episode $i$, the cost module updates $\gamma_{m,i+1}$ using (21) and (22):
$\gamma_{m,i+1}=\frac{N_{i}(\bm{s}_{0},\bm{a}^{\textsc{h}}_{m,i}(T))}{D_{i}(\bm{s}_{0},\bm{a}^{\textsc{h}}_{m,i}(T))},~{}~{}i\in[E],$ (24)
where $T$ is the stepped needed set to be the same for every episode, as motivated in Appendix B. Eq. (24) corresponds to the update procedure of the quotient coefficient in (19) as in the outer loop of the fractional RL framework.
(a)            (b)
Figure 3: Convergence of (a) average AoI and (b) the average value of quotient coefficients $\gamma_{m,i}$ across devices, where $\gamma_{m,i}$ is updated every $50$ episodes.
(a)            (b)          (c)          (d)
Figure 4: Performance under different (a) processing capacities of edges, (b) drop coefficients, (c) task densities, and (d) processing capacities of mobile devices.
6 Performance Evaluation
We perform experiments to evaluate our proposed fractional DRL algorithm. We consider two edge nodes and 20 mobile devices learning their own policies simultaneously. Unless otherwise specified, we follow the experimental settings in (Tang and Wong 2022, Table I). We present more detailed experiment settings in Appendix D.
We denote our proposed approach with Frac. DRL (OFL+U), which is short for “fractional DRL with offloading and updating policies”. This is compared with several benchmark methods:
  • Random scheduling: The updating and offloading decisions are randomly generated within action space.
  • PGOA (Yang et al. 2018): This corresponds to a best response algorithm for potential game in MEC systems.
  • Non-Fractional DRL (denoted by Non-Frac. DRL) (Xu et al. 2022): This benchmark adopts D3QN network to learn the offloading policy. In contrast to our proposed framework, this benchmark is non-fractional. That is, its objective approximates the ratio-of-expectation average AoI in (5) by an expectation-of-ratio expression: $\mathop{\text{minimize}}_{\bm{\pi}_{m}}~{}~{}\mathbb{E}\left[\left.\frac{A(Y_{m,k},Z_{m,k+1},Y_{m,k+1})}{Y_{m,k}+Z_{m,k+1}}\right|\bm{\pi}_{m}\right].$ Such an approximation can circumvent the fractional challenge but incurs large accuracy loss.
  • Frac. DRL with Offloading Only (denoted by Frac. DRL (OFL)): We propose this algorithm by simplifying our fractional DRL algorithm through considering only the offloading policy. The updating policy (i.e., $Z_{m,k}$) is set to zero, as in many of the existing works (Xie, Wang, and Weng 2022; He et al. 2022; Xu et al. 2022).
The performance difference between Non-Frac. DRL and Frac. DRL (OFL) shows the significance of our proposed fractional DRL. The difference between Frac. DRL (OFL) and Frac. DRL (OFL+U) shows the necessity of joint offloading and waiting optimization.
Convergence: Fig. 3 illustrates the convergence of our proposed Frac. DRL (OFL) and Frac. DRL (OFL+U) algorithms. Unlike non-fractional approaches, our proposed approach involves the convergence of not only the neural network (see Fig. 3(a)) but also the quotient coefficient $\gamma$ (see Fig. 3(b)). As a result, the convergence curve of AoI may sometimes change non-monotonically. In Fig. 3(a), both Frac. DRL (OFL) and Frac. DRL (OFL+U) converge after roughly 350 episodes. As for the converged AoI, Frac. DRL (OFL+U) outperforms Frac. DRL (OFL) by 31.3%.
Edge Capacity: Fig. 4(a) evaluates the performance of our proposed schemes under different node processing capacities. First, the proposed fractional DRL-based algorithm consistently achieves lower average AoI, compared against those non-fractional benchmarks. Such an advantage is more significant as the per-node processing capacity is larger. When processing capacity is $75$ GHz, Frac. DRL (OFL+U) can achieve an average AoI reduction of $54.8\%$ and $73.6\%$, compared against PGOA and Non-Frac. DRL, respectively. Second, when compared with Frac. DRL (OFL), Frac. DRL (OFL+U) can further reduce the average AoI up to $19.9\%$ when the processing capacity of edge nodes is 55 GHz. This further shows that, in addition to the fractional RL framework, a well-designed updating policy also plays an important role of further improving the performance, especially there are relatively high edge loads.
Drop Coefficient: In Fig. 4(b), we consider different drop coefficients, i.e., the ratio of the drop time $\bar{Y}$ to the average time of processing a task. The performance gaps between the proposed schemes and benchmarks are large when the drop ratio gets smal (i.e., tasks are more delay-sensitive). When the drop coefficient is 1.6, the Frac. DRL (OFL+U) reduces the average AoI by $55.1\%$, compared with Non-Frac. DRL.
Task Density: In Fig. 4(c), we evaluate algorithm performance under different task densities, which affect the expected processing time of tasks at both edge nodes and mobile devices. Specifically, our proposed Frac. DRL (OFL) and Frac. DRL (OFL+U) schemes outperform all the benchmarks. In addition, the performance gaps increase as the task density increases, which shows the benefit of our proposed algorithm under large task densities. When the task density is 0.65, Frac. DRL (OFL+U) achieves an average reductions of $38.4\%$ and $49.3\%$, compared against PGOA and Non-Frac. DRL, respectively. Meanwhile, Frac. DRL (OFL+U) outperforms Frac. DRL (OFL) by up to $24.2\%$.
Mobile Capacity: In Fig. 4(d), as the processing capacity of mobile devices decreases, the gap between Frac. DRL (OFL) and Non-Frac. DRL significantly increases, indicating the necessity of our fractional scheme. When mobile capacity is 2 GHz, Frac. DRL (OFL+U) can achieve an average AoI reduction of 48.8% compared to PGOA.
To summarize, our proposed schemes significantly outperform non-fractional benchmarks, especially under large task density, delay-sensitive tasks, and small mobile device processing capacity. Meanwhile, the joint optimization over offloading and updating can further increase the system performance by up to $31.3$%. We present additional convergence and performance evaluation under different networks hyperparamters, distribution of processing duration and scale of mobile devices in Appendix D.
7 Conclusion
This paper has studied the computational task scheduling (including offloading and updating) problem for age-minimal MEC. To address the underlying challenges of unknown load dynamics and the fractional objective, we have proposed a fractional RL framework with a provable linear convergence rate. We further designed a fractional DRL algorithm that incorporates D3QN and DDPG techniques to tackle hybrid action space. Experimental results show that our proposed fractional algorithms significantly reduce the average AoI, compared against several benchmarks. Meanwhile, the joint optimization of offloading and updating can further reduce the average AoI, validating the effectiveness of our proposed scheme. There are several future directions, including incorporating RL techniques for non-stationary environments and social optimal scheduling.
Appendix A: Proof of Theorem 1
Define
$\bm{a}^{*}(\gamma)$ $\triangleq\arg\max_{a\in\mathcal{A}}\left(N_{\gamma}(\bm{s}_{0},\bm{a})-\gamma D_{\gamma}(\bm{s}_{0},\bm{a})\right),$ (25a)
$Q({\gamma})$ $\triangleq\max_{\bm{a}\in\mathcal{A}}\left(N_{\gamma}(\bm{s}_{0},\bm{a})-\gamma D_{\gamma}(\bm{s}_{0},\bm{a})\right),$ (25b)
$N(\gamma)$ $=N_{\gamma}(s_{0},\bm{a}^{*}(\gamma))~{}~{}{\rm and}~{}~{}D(\gamma)=D_{\gamma}(s_{0},\bm{a}^{*}(\gamma)),$ (25c)
$F(\gamma)$ $=\frac{N(\gamma)}{D(\gamma)},$ (25d)
$\bm{a}_{i}$ $\triangleq\arg\max_{\bm{a}\in\mathcal{A}}\left(N_{k}(\bm{s}_{0},\bm{a})-\gamma D_{k}(\bm{s}_{0},\bm{a})\right),$ (25e)
for all $\gamma\geq 0$. In the remaining part of this proof, we use $Q_{i}=Q_{i}(\bm{s}_{0},\bm{a}_{i})$, $N_{i}=N_{i}(\bm{s}_{0},\bm{a}_{i})$, and $D_{i}=D_{i}(\bm{s}_{0},\bm{a}_{i})$ for presentation simplicity. Note that
$\frac{N(\gamma^{\prime})}{D(\gamma^{\prime})}-\frac{N_{i}}{D_{i}}$ (26)
$\overset{(a)}{\geq}~{}$ $\frac{N(\gamma^{\prime})}{D(\gamma^{\prime})}-\frac{N(\gamma^{\prime})}{D_{i}}-\gamma\left[\frac{D(\gamma^{\prime})}{D_{i}}-\frac{D(\gamma^{\prime})}{D(\gamma)}\right]-\frac{\epsilon_{i}}{D_{i}}$
[…]
where (a) is from the suboptimality of $N_{i}$. Sequences $\{\gamma_{i}\}$, $\{Q_{i}\}$, and $\{D_{i}\}$ generated by FQL Algorithm satisfy $\gamma_{i}-\gamma^{*}\geq\gamma_{i}-{N_{i}}/{D_{k}}=-{Q_{i}}/{D_{i}}\geq{\epsilon_{i}}/{\alpha D_{i}}$ for all $i$ such that $Q_{i}<0$. In addition, from the fact that $N_{i}-\gamma_{i}D_{i}\leq N(\gamma^{*})-\gamma_{i}D(\gamma^{*})$ and $N(\gamma^{*})-\gamma^{*}D(\gamma^{*})\leq N_{i}-\gamma^{*}D_{i}$ for all $i\in[E]$, it follows that $(\gamma_{i}-\gamma^{*})(D(\gamma^{*})-D_{i})\leq 0,~{}\forall i\in[E],$ and hence $\gamma_{i}\geq\gamma^{*}~{}~{}{\rm if~{}and~{}only~{}if}~{}D_{i}\geq D(\gamma^{*}).$ It follows from (26) that, for all $i$,
$\gamma_{i+1}-\gamma^{*}=F_{i}-\gamma^{*}\leq~{}~{}$ $(\gamma_{i}-\gamma^{*})\left(1-\frac{D(\gamma^{*})}{D_{i}}\right)+\frac{\epsilon_{i}}{D_{i}}$
$\overset{(b)}{\leq}~{}~{}$ $(\gamma_{i}-\gamma^{*})\left(1+\alpha-\frac{D(\gamma^{*})}{D_{i}}\right).$
Note that $(b)$ involves induction. Specifically, if $\gamma_{i}<\gamma^{*}$, then $\gamma_{i+1}<\gamma^{*}$. Therefore, we have that, if $Q_{i}<0$ then $Q_{i+1}<0$, and hence that $\epsilon_{i}\leq-Q_{i}$ for all $i\in[E]$. Since $\alpha\in(0,1)$, it follows that $(\gamma_{i+1}-\gamma^{*})/(\gamma_{i}-\gamma^{*})\in(0,1)$ for all large enough $i$, which implies that $\{\gamma_{i}\}$ converges linearly to $\gamma^{*}$.
Appendix B: Analysis of Time Complexity
Proposition 1.
Proposed FQL Algorithm satisfies the stopping condition $\epsilon_{i}<-\alpha Q_{i}(\bm{s}_{0},\bm{a}_{i})$ for some $\alpha\in(0,1)$, then after
$T_{i}=\left\lceil\frac{11.66\log(2|\mathcal{Z}|/(E\zeta))}{\alpha^{2}}\right\rceil$ (27)
steps of SQL, the uniform approximation error $\lVert Q^{*}_{\gamma_{i}}-Q_{i}\rVert\leq\epsilon_{i}$ holds for all $i\in[E]$, with a probability of $1-\zeta$ for any $\zeta\in(0,1)$.
[…]
Proposition 1 shows that the total steps needed $T_{i}$ does not increase in $i$, even though the stopping condition $\epsilon_{i}<-\alpha Q_{i}(\bm{s}_{0},\bm{a}_{i})$ is getting more restrictive as $i$ increases.
Appendix C: DRL Network
DDPG Module
[…]
Action Selection
[…]
$\bm{a}^{\textsc{u}}_{m}(k)=\mu_{m}(\bm{s}^{\textsc{u}}_{m}(k)|\bm{\theta}^{\mu}_{m})+\mathcal{N}_{m},$ (28)
where $\mathcal{N}_{m}$ is an exploration noise, and $\mu_{m}(\bm{s}^{\textsc{u}}_{m}(k)|\bm{\theta}^{\mu}_{m})$ denotes the action policy of state $\bm{s}^{\textsc{u}}_{m}(k)$ with $\bm{\theta}^{\mu}_{m}$.
Neural Network Training
[…]
$L=\frac{1}{|\mathcal{K}_{b}|}\sum_{k\in\mathcal{K}_{b}}\left(y_{m,k}-Q^{\textsc{u}}_{m}(\bm{s}^{\textsc{u}}_{m}(k),\bm{a}^{\textsc{u}}_{m}(k)|\bm{\theta}^{\textsc{u}}_{m})\right)^{2},$ (29)
and $y_{m,k}\!=\!c_{m}(k)\!+\!\delta Q^{\textsc{u}}_{m}\!(\bm{s}^{\textsc{u}}_{m}(k\!+\!1),\mu_{m}\!(\bm{s}^{\textsc{u}}_{m}(k+\!1)|\bm{\theta}^{\mu-}_{m})\!\\ |\bm{\theta}^{\textsc{u}-}_{m})$. In addition, the DDPG module updates the actor policy $\bm{\theta}^{\mu}_{m}$ using the sampled policy gradient: for all $m\in\mathcal{M}$,
$\nabla_{\bm{\theta}^{\mu}_{m}}J\approx\frac{1}{|\mathcal{K}_{b}|}\sum_{k\in\mathcal{K}_{b}}$ $\nabla_{\bm{a}^{\textsc{u}}_{m}}Q^{\textsc{u}}_{m}(\bm{s}^{\textsc{u}}_{m}(k),\mu_{m}(\bm{s}^{\textsc{u}}_{m}(k)|\bm{\theta}^{\mu}_{m})|\bm{\theta}^{\textsc{u}}_{m})$
×$\times\nabla_{\bm{\theta}^{\textsc{u}}_{m}}\mu_{m}(\bm{s}^{\textsc{u}}_{m}(k)|\bm{\theta}^{\mu}_{m}).$ (30)
Finally, the DDPG module uses soft target updates, based on $\bm{\theta}^{\textsc{u}-}_{m}=\tau\bm{\theta}^{\textsc{u}}_{m}+(1-\tau)\bm{\theta}^{\textsc{u}-}_{m}$ and $\bm{\theta}^{\mu-}_{m}=\tau\bm{\theta}^{\mu}_{m}+(1-\tau)\bm{\theta}^{\mu-}_{m}$, with a small $\tau$.
D3QN Module
[…]
Action Selection
[…]
$\bm{a}^{\textsc{o}}_{m}(k)=\left\{\begin{array}[]{ll}\text{a random action from $\mathcal{A}^{\textsc{o}}$},&\text{w.p.}\quad\epsilon_{r},\\ \arg\min_{\bm{a}}Q^{\textsc{o}}_{m}(\bm{s}^{\textsc{o}}_{m}(k),\bm{a};\bm{\theta}^{\textsc{o}}_{m}),&\text{w.p.}\quad 1-\epsilon_{r},\end{array}\right.$ (31)
where ‘w.p.’ refers to “with probability”, and $\epsilon_{r}$ is the probability of random exploration.
Neural Network Training
When task $k$ of mobile device $m$ has been processed, D3QN module observes the cost $c_{m}(k)$ and stores experience $(\bm{a}^{\textsc{o}}_{m}(k),\bm{s}^{\textsc{o}}_{m}(k),c_{m}(k),\bm{s}^{\textsc{o}}_{m}(k+1))$ to replay buffer. Then, the D3QN module randomly samples mini-batches to update $\bm{\theta}^{\textsc{o}}_{m}$ by minimizing the difference between the recent Q-value of the selected action under the observed state and the target Q-value $\bm{Q^{\textsc{trg}}}_{m}=\{Q^{\textsc{trg}}_{m,k}\}_{k\in\mathcal{K}_{b}}$:
$L(\bm{\theta}^{\textsc{o}}_{m},\bm{Q^{\textsc{trg}}}_{m})=\frac{1}{|\mathcal{K}_{b}|}\sum_{k\in\mathcal{K}_{b}}\Big{(}Q^{\textsc{o}}_{m}(\bm{s}^{\textsc{o}}_{m}(k),\bm{a}_{m}(k);\bm{\theta}^{\textsc{o}}_{m})-Q^{\textsc{trg}}_{m,k}\Big{)}^{2}.$ (32)
[…]
$Q^{\textsc{trg}}_{m,k}=c_{m}(k)+\delta Q^{\textsc{o}}_{m}(\bm{s}_{m}(k+1),\bm{a}^{\text{Next}}_{m,k};\bm{\theta}^{\textsc{o}-}_{m}),$ (33)
[…]
$\bm{a}^{\text{Next}}_{m,k}=\arg\min_{\bm{a}\in\mathcal{A}^{\textsc{o}}}Q^{\textsc{o}}_{m}(\bm{s}_{m}(k+1),\bm{a};\bm{\theta}_{m}^{o}).$ (34)
Hyperparameter of Neural Network
For the D3QN networks, we use RMSProp optimizer. The batch size is $32$, the learning rate is $3\times 10^{-4}$, and the discount factor is $\delta=0.9$. The probability of random exploration $\epsilon_{r}$ in (31) is gradually decreasing from $1$ to $0.003$. For the DDPG networks, we use Adam as the optimizer. The batch size is $64$, and the learning rates are $1\times 10^{-4}$ and $1\times 10^{-3}$ for actor and critic networks, respectively. Detailed structures of above networks see the codes provided.
Appendix D: Additional Evaluation
We evaluate the convergence and performance of our algorithms Frac. DRL (OFL) and Frac. DRL (OFL+U) under different hyperparameters and environment settings.
Environment Setting
The proposed DRL framework is trained online with an infinite-horizon continuous-time environment, where we train the D3QN and DDPG networks to upgrade the task updating and offloading decisions respectively with collected experience. We evaluate the convergence of variants of our proposed DRL framework Frac. DRL (OFL) and Frac. DRL (OFL+U) under various DRL network hyperparameters respectively. Basically, we consider 1000 episodes (1500 episodes if necessary) with constant time limit and update fractional coefficient $\gamma_{m,i}$ every 50 episodes and for every experiment point, we run three times and average the results in our evaluation. We develop our programme on AMD EPYC 7402 CPU and Nvidia RTX 3090Ti GPU with Utuntu 20.04. Detailed package versions are presented in our Code Appendix. We present default basic parameter settings of our MEC environment in Table 1.
Table 1: Environment parameter settings
Parameter Value
Number of mobile devices 20
Number of edge nodes 2
Capacity of mobile devices 2.5 GHz
Capacity of edge nodes 41.8 GHz
Task size 30 Mbits
Task density 0.297 gigacycles per Mbits
Drop coefficient 1.5
Hyperparameter Evaluation
First we evaluate the convergence of the variant Frac. DRL (OFL) in Figure 5, which has offloading decisions with D3QN networks only and we keep hyperparamters to be the same among same type of networks from different mobile devices. In the figures, the $x$-axis represents the training episode, and the $y$-axis shows the averaged AoI among the mobile devices in each episode. The performance evaluations are plotted under different hyperparameters of the neural networks and we denote the random scheduling policy “Random”, where we randomly choose the offloading actions.
Each mobile device performs the proposed algorithm in a decentralized manner without interacting with other mobile devices. Note that even under such an independent training framework, the scheduling policy of each mobile device can gradually improve and converge. This algorithm contains three modules: cost module, DDPG module, and D3QN module. The cost module determines the cost function for the DDPG and the D3QN modules based on the proposed fractional RL framework. The DDPG and the D3QN modules are responsible for making the task updating and offloading decisions, respectively.
Apart from these decisions made by neural networks, we have an additional dropping scheme in our algorithm. When the task processing duration of a mobile device exceeds the limitation, the task is dropped and start a new task immediately meanwhile the recorded AoI of the mobile device keeps increasing. This scheme can significantly better the performance.
(a)            (b) Refer to caption
Figure 5: Frac. DRL (OFL) convergence and performance under different (a) learning rates, (b) batch sizes, (c) optimizers.
Fig.5(a) shows the convergence under different values of learning rates (denoted “lr”) of D3QN networks, which is used to scale the magnitude of parameter updates during gradient descent. As shown in Fig.5(a), lr = $3\times 10^{-4}$ leads to a relatively smooth convergence and small average AoI. If the learning rate is too large (i.e., $3\times 10^{-1}$), it will be hard to converge and if it is too small (i.e., $3\times 10^{-5}$), it converges slowly. Fig.5(b) shows Frac. DRL (OFL) performance under different batch sizes, i.e., the number of samples that will be propagated through the network. We can see when batch size is 32, the algorithm results in a promising performance and the result gets worse when batch size is too small (i.e. batch size = 8). Fig.5(c) shows Frac. DRL (OFL) performance under different optimizers, which consist of adaptive moment estimation (Adam), gradient descent (denoted by ”GD”) and RMSprop optimizers, which are methods used to update the neural network to reduce the losses. In Fig.5(c), RMSProp and Adam optimizers achieve similar convergent average AoI which is far better than gradient descent.
(a)            (b) Refer to caption
Figure 6: Frac. DRL (OFL+U) convergence and performance under different (a) learning rates of Actor and Critic Networks respectively, (b) batch sizes, (c) optimizers.
Then, we keep the hyperparameters of D3QN networks be constant and experiment the convergence and performance of Frac. DRL (OFL+U) algorithm under different hyperparameters of DDPG networks which are kept the same among different mobile devices. Fig.6(a) shows the convergence under different values of learning rates of DDPG networks. As shown in Fig.5(a), the learning rates of Actor and Critic being $3\times 10^{-4}$ and $3\times 10^{-3}$ respectively leads to a relatively fast convergence and small convergent average AoI. If the learning rates get larger, it results in worse convergence and performance. Fig.6(b) shows Frac. DRL (OFL) performance under different batch sizes. We can see the convergent average AoI when batch size is 64 is better than the performance when batch sizes increase or decrease. Fig.5(c) shows Frac. DRL (OFL) performance under different optimizers, where all three optimizers have similar performance.
(a)            (b)
Figure 7: Convergence under exponential and lognormal distributions of (a) of Frac. DRL (OFL), (b) Frac. DRL (OFL+U).
Figure 8: Performance comparison under different numbers of mobile devices.
Distributions of Duration
We experiment our algorithms under widely used exponential distribution and lognormal distribution, which is suitable for communication latency. In Fig.8, we show the convergence of our proposed algorithms Frac. DRL (OFL) and Frac. DRL (OFL+U) under these two distributions, where the exponential and lognormal distributions are denoted by ”Exp.” and ”Log” respectively. It shows that both our proposed algorithms can converge with similar speed under these two distributions.
Number of Mobile Devices
In Fig.7, we compare the algorithm performance under different numbers of mobile devices. The performance gaps between the proposed fractional approaches and the non-fractional benchmarks are larger when the number mobile device is small. When the number of mobile devices is equal to 10, the Frac. DRL (OFL+U) reduces the average AoI by 68.9% when compared with Non-Frac. DRL.
9/sVg8PXweDAauqqr2cDjEer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">